# FedVSR: Towards Model-Agnostic Federated Learning in Video Super-Resolution

### v1

@@ Abstract
Video Super-Resolution (VSR) reconstructs high-resolution videos from low-resolution inputs to restore fine details and improve visual clarity. While deep learning-based VSR methods achieve impressive results, their centralized nature raises serious privacy concerns, particularly in applications with strict privacy requirements. Federated Learning (FL) offers an alternative approach, but existing FL methods struggle with low-level vision tasks, leading to suboptimal reconstructions. To address this, we propose FedVSR [^1] Preprint. Under review., a novel, architecture-independent, and stateless FL framework for VSR. Our approach introduces a lightweight loss term that improves local optimization and guides global aggregation with minimal computational overhead. To the best of our knowledge, this is the first attempt at federated VSR. Extensive experiments show that FedVSR outperforms general FL methods by an average of 0.85 dB in PSNR, highlighting its effectiveness. The code is available at: https://github.com/alimd94/FedVSR
@@ 1 Introduction
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[31] Refer to caption (d) FedProx[23]
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[18] Refer to caption (g) FedVSR (Ours)
@@
Figure 1: Comparison of RVRT[24] output trained with different Federated Learning algorithms. General FL methods struggle to reconstruct fine details and textures.
@@
Video Super-Resolution (VSR) is an advanced computational process that reconstructs high-resolution (HR) video sequences from low-resolution (LR) counterparts leveraging both spatial and temporal information [27]. Unlike Single Image Super-Resolution (SISR), which relies solely on intra-frame details, VSR exploits temporal correlations across consecutive frames to enhance the quality and sharpness of the output. This technique is particularly valuable in real-world applications such as surveillance [10], video streaming [20], and medical imaging [40, 37], where improving visual fidelity and maintaining temporal coherence are critical. VSR addresses complex and diverse degradations commonly found in low-quality videos, including down-sampling [54], noise [57], blur [33], and compression artifacts [9], while ensuring the perceptual quality and consistency of the restored video. However, the data-driven and centralized nature of existing deep learning-based VSR methods introduces a potential risk of privacy breaches, particularly when handling sensitive user data that require strict confidentiality.
Federated Learning (FL) presents a compelling alternative to mitigate privacy risks associated with centralized training in Deep Learning-based VSR methods. As a privacy-preserving framework, FL allows multiple participants or edge devices to collaboratively train a shared Machine Learning model without sharing raw data, effectively addressing data security concerns [31]. Unlike conventional distributed methods that rely on frequent synchronization, FL enables selected participants to perform multiple local updates before their models are aggregated into a global one. This approach minimizes communication overhead and adapts to the limited computational resources of edge devices. By avoiding centralized data collection, FL ensures the protection of sensitive information, offering a practical solution for privacy-aware applications in video super-resolution and similar domains.
Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as Image Classification [18, 23, 46], Face Recognition [36, 29], and Segmentation [32, 29]. However, their application in Super-Resolution (SR) remains largely unexplored. Studies such as [34, 51] indicate that directly applying existing FL techniques to SISR results in poor-quality reconstructions with blurred textures. This limitation arises from their inability to effectively capture rich low-level information. Motivated by this observation, we propose an FL-based approach for VSR that aims to generate better quality output while ensuring data privacy.
In this paper, we propose FedVSR a novel model-agnostic framework for VSR under the FL paradigm, which treats the VSR models as a black-box, making it highly adaptable to different VSR methods and future advancements. To the best of our knowledge, this is the first attempt to address the VSR problem within a decentralized setting. Moreover, our framework does not retain previous model weights for aggregation. Instead, we introduce an effective loss term based on Discrete Wavelet Transform (DWT) which not only enhances local training efficiency with negligible computational overhead but also informs the global aggregation process, leading to improved reconstruction quality. This design enables our approach to efficiently capture essential information while ensuring strict data privacy, making it well-suited for real-world applications where both visual fidelity and confidentiality are critical. In summary, the main contributions of this paper are as follows:
  • We propose a DWT-based loss term to enhance high-frequency detail in Federated VSR. This term also serves as a reliable performance metric for weighting parameter aggregation.
  • Our approach is fully model-agnostic and stateless, requiring no architecture-specific assumptions or storage of past weights, making it practical and scalable.
@@
  • Extensive experiments validate our method’s effectiveness, outperforming existing approaches and paving the way for privacy-preserving low-level vision in a federated setting.
  • To the best of our knowledge, this is the first work addressing federated VSR, balancing privacy preservation with improved visual results.
@@ 3.1 VSR Problem Definition
In addition to the widely used RGB color space, other representations such as the YUV and YCbCr color spaces are frequently employed in VSR applications. Let $F_{i}\in\mathbb{R}^{H\times W\times 3}$ denote the $i$-th frame in a given LR video sequence, while $\hat{F}_{i}\in\mathbb{R}^{sH\times sW\times 3}$ represents its corresponding high-resolution (HR) counterpart, where $s$ is the scaling factor (e.g., $s=2,4,8$). Additionally, we define a set of $12N+$ 1 HR frames $\{\hat{I}_{j}\}_{j=i-N}^{i+N}$ centered around $\hat{I}_{i}$, where $N$ represents the temporal radius. The degradation process that generates an LR video sequence from its HR counterpart can be mathematically formulated as: {scriptsize }
@@
A more detailed expression for the degradation process can be written as: {scriptsize }
@@
where $D$ represents the down-sampling operation, $B$ denotes the blurring effect, $E_{i\to j}$ is a warping function that accounts for motion between frames $\hat{I}_{i}$ and $\hat{I}_{j}$, and $n_{j}$ represents the image noise. The goal of a VSR algorithm is to reconstruct an HR video sequence from degraded LR frames in a manner that closely resembles the ground truth video. This makes this task an ill-posed problem because there exist infinitely many possible HR images $\hat{I}_{i}$ that can produce the same LR image $I_{i}$ after degrading, meaning the degradation operator is highly non-injective [44]. While some of the techniques used in ISIR can be applied to individual frames, VSR must also consider inter-frame dependencies to ensure motion coherence across the sequence. The inverse process of degradation, which represents the super-resolution task, can be expressed as: {scriptsize }
@@ 3.2 FL Problem Definition
FL aims to minimize the following objective function across participating clients: {scriptsize }
@@
Each client updates the global model locally: {scriptsize }
@@
where $\eta$ is the learning rate. The updated models are aggregated by the central server using a weighted averaging scheme: {scriptsize }
@@ 4.1.1 DWT for High-Frequency Preservation
The coefficient in those four sub-bands are calculated as: {scriptsize }
@@
{scriptsize }
$L_{i}^{LH}=\bar{F}_{i}^{LH}-\hat{F}_{i}^{LH},\quad L_{i}^{HL}=\bar{F}_{i}^{HL}-\hat{F}_{i}^{HL},\quad L_{i}^{HH}=\bar{F}_{i}^{HH}-\hat{F}_{i}^{HH}$ (9)
Then with the Charbonnier loss [7] we calculate the helper term $\mathcal{L}_{\text{DWT}}$ : {scriptsize }
@@
Now, we define the overall loss for the model with two weighted terms, one for the generic VSR loss $\mathcal{L}_{\text{VSR}}$ and one for the DWT-based high-frequency loss $\mathcal{L}_{\text{HiFr}}$ as follows: {scriptsize }
@@ 5 Experiments
Model Frames, Params(M) Algorithm Less Heterogeneous More Heterogeneous
@@
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
VRT[25] 16, 35.6 FedAvg[31] 29.71 0.8522 24.88 0.7680 28.85 0.8282 24.60 0.7516
FedProx[23] 28.82 0.8281 24.18 0.7341 28.64 0.8236 23.99 0.7256
SCAFFOLD[18] 28.59 0.8181 23.99 0.7182 28.17 0.8074 24.04 0.7208
FedVSR (Ours) 29.86 0.8562 24.92 0.7682 29.03 0.8339 24.73 0.7584
RVRT[24] 30, 10.8 FedAvg[31] 31.30 0.8877 25.96 0.8156 29.89 0.8578 25.67 0.8040
FedProx[23] 30.09 0.8618 25.20 0.7775 29.59 0.8526 25.02 0.7730
SCAFFOLD[18] 29.63 0.7967 24.79 0.7317 26.71 0.7142 23.08 0.6441
FedVSR (Ours) 31.49 0.8908 26.05 0.8191 30.34 0.8678 25.90 0.8048
IART[49] 16, 13.4 FedAvg[31] 30.56 0.8729 25.43 0.7949 29.52 0.8493 25.04 0.7774
FedProx[23] 29.63 0.8500 24.72 0.7599 29.30 0.8444 24.42 0.7505
SCAFFOLD[18] 30.29 0.8678 25.18 0.7857 29.26 0.8413 24.85 0.7690
FedVSR (Ours) 30.88 0.8812 25.54 0.8012 29.70 0.8518 25.26 0.7850
Table 1: PSNR and SSIM comparison of FL algorithms across different VSR models and datasets under varying levels of heterogeneity.
@@
For our baseline, we use a combination of FedAvg[31], FedProx[23], and SCAFFOLD[18] from the FL side, along with VRT[25], RVRT[24], and IART[49] from VSR.
@@
The quality of the reconstructed video is typically evaluated using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM).
@@ 5.2 Results and Comparison
As shown in Table 1, our proposed FedVSR algorithm consistently outperforms other federated learning (FL) algorithms across all tested models and datasets. On average, FedVSR surpasses existing FL methods by a margin of 0.89 dB in PSNR and 0.028 in SSIM in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.80 dB in PSNR and 0.027 in SSIM, demonstrating its robustness under challenging conditions. Among the evaluated models, RVRT benefits the most from FedVSR, achieving the highest performance margin. In contrast, IART shows the lowest overall improvement but still maintains a competitive edge over the baseline FL methods. When comparing FedVSR only against the second-best FL algorithm, FedAvg, our approach achieves an average PSNR margin of 0.22 dB in less heterogeneous settings and 0.27 dB in more heterogeneous settings. Similarly, the SSIM improvement over FedAvg is 0.005 in less heterogeneous cases and 0.006 in more heterogeneous cases. The smallest improvement relative to FedAvg is observed for VRT in PSNR and IART in SSIM, while the highest improvement is consistently achieved by RVRT, particularly in more heterogeneous settings, further validating FedVSR’s ability to handle real-world federated VSR challenges effectively. The findings emphasize the importance of high-frequency-aware optimization in VSR FL-based training.
The observed performance trends in our experiments align with existing findings on FedProx and SCAFFOLD, further validating our results. As shown in Fig.2, FedVSR consistently achieves superior PSNR values across different VSR models and datasets, outperforming FedAvg, FedProx, and SCAFFOLD, particularly in later training rounds. This aligns with prior research on FedProx, which highlights its failure to reach the stationary points of the global optimization objective, even under homogeneous data distributions. The study suggests that FedProx’s limiting points may remain far from the true minimizer of the global empirical risk [43], which could explain its suboptimal performance in our experiments. Although FedProx may still generalize well by aggregating local data effectively in certain conditions, its weaker convergence properties become apparent in more heterogeneous scenarios. We also observe that FedProx is highly sensitive to the $\mu$ value during training. The original paper suggests $\{0.001,0.01,0.1,1\}$ as candidate values and finds that $\mu=1$ performs better in most cases. However, we notice that for any $\mu$ greater than $31e-$ 3, the VSR model struggles significantly with learning. This suggests that the relationship is more complex and requires further investigation and analysis.
@@
As showen in Fig. 1 clearly demonstrate the limitations of FedAvg, FedProx, and SCAFFOLD in accurately reconstructing fine details. While these methods reduce some pixelation from the LQ input, they all fail to recover the textures and sharp edges present in the ground-truth. FedAvg produces overly smooth results, blurring important structural details such as the wooden lattice and frame edges. FedProx exhibits similar blurring, struggling to restore high-frequency textures. SCAFFOLD, not only fails to produce a sharp reconstruction but also introduces noticeable artifacts, distorting the texture of the wooden lattice and making the output appear unnatural. In contrast, FedVSR delivers significantly sharper and more realistic reconstructions, clearly outperforming the other approaches. Due to space limitations, additional qualitative and quantitative results are provided in the supplementary material 7.

### v2

@@ Abstract
Video super-resolution aims to enhance low-resolution videos by leveraging both spatial and temporal information. While deep learning has led to impressive progress, it typically requires centralized data, which raises privacy concerns. Federated learning offers a privacy-friendly solution, but general FL frameworks often struggle with low-level vision tasks, resulting in blurry, low-quality outputs. To address this, we introduce FedVSR, the first FL framework specifically designed for VSR. It is model-agnostic and stateless, and introduces a lightweight loss function based on the DWT to better preserve high-frequency details during local training. Additionally, a loss-aware aggregation strategy combines both DWT-based and task-specific losses to guide global updates effectively. Extensive experiments across multiple VSR models and datasets demonstrate that FedVSR consistently outperforms existing FL methods, achieving up to 0.82 dB higher PSNR, 0.0327 higher SSIM, and 0.0251 lower LPIPS. These results underscore FedVSR’s ability to bridge the gap between privacy and performance, setting a new benchmark for federated learning in low-level vision tasks. The code is available at: https://github.com/alimd94/FedVSR
@@ 1 Introduction
Refer to caption (b) LQ ($\times$4) Refer to caption (c) FedAvg[36] 30.41, 0.8600, 0.1777 Refer to caption (d) FedProx[27] 30.07, 0.8531, 0.1851
Refer to caption (e) GT Refer to caption (f) SCAFFOLD[21] 26.78, 0.7148, 0.2516 Refer to caption (g) FedVSR (Ours) 32.00, 0.8889, 0.1508
@@
Figure 1: Comparison of RVRT[29] output trained with different federated learning algorithms. General FL methods struggle to reconstruct fine details and textures. PSNR, SSIM, and LPIPS metrics are shown in red, blue, and green respectively.
@@
Video super-resolution (VSR) is an advanced computational process that reconstructs high-resolution (HR) video sequences from low-resolution (LR) counterparts, leveraging both spatial and temporal information [32]. Unlike single-image super-resolution (SISR), which relies solely on intra-frame details, VSR exploits temporal correlations across consecutive frames to enhance the quality and sharpness of the output. This capability is crucial in real-world applications such as surveillance [12], video streaming [23], and medical imaging [45, 42], where visual fidelity and temporal coherence are mission-critical. In practice, VSR must also cope with degradations such as down-sampling [60], noise [63], blur [38], and compression artifacts [11]. However, modern deep-learning pipelines address these degradations with data-hungry networks trained on centrally collected data, but centralization itself raises privacy, regulatory, and bandwidth concerns, especially when the raw videos contain sensitive or personally identifiable information.
Federated learning (FL) provides a principled remedy to these concerns [36]. In many VSR scenarios, such as smartphones, surveillance, health-related domains, and remote edge sensors, videos are captured on devices with limited bandwidth and strict privacy constraints. Uploading hours of footage can consume tens of gigabytes of storage; for instance, even a short 10-second 8K video at 30 fps in raw format can exceed 10 GB in size [2]. Additionally, uploading such data may violate policies and laws such as GDPR or HIPAA [19]. Unlike conventional distributed methods that rely on frequent synchronization, FL enables selected participants to perform multiple local updates before their models are aggregated into a global one. This approach minimizes communication overhead and accommodates the limited computational resources of edge devices. These combined benefits make FL a logical and effective foundation for VSR systems that face such concerns.
Despite significant advancements in FL methods, most research has focused on high-level or downstream visual tasks such as image classification [21, 27, 51], facial recognition [41, 34], and segmentation [37, 34]. However, their application in super-resolution (SR) remains largely unexplored. Existing studies [39, 56] found that directly applying existing general FL techniques to SR tasks results in poor-quality reconstructions with blurred textures, mainly due to their inability to effectively capture rich low-level information such as texture, edges and motion. Motivated by these limitations, we propose FedVSR, a novel model-agnostic framework for VSR under FL. FedVSR treats VSR models as black boxes, allowing it to work with a variety of architectures without requiring architectural changes or access to intermediate layers. To improve performance under FL constraints, we introduce a lightweight loss term based on the Discrete Wavelet Transform (DWT), which emphasizes high-frequency details that are often lost during decentralized training. While high-frequency preservation is valuable for all SR tasks, its importance is amplified in VSR due to the added complexity of maintaining temporal coherence across frames. In federated settings, data heterogeneity and limited communication intensify this challenge. FedVSR addresses these issues without maintaining state between rounds, making it both scalable and practical for privacy-sensitive, resource-constrained environments. In summary, the main contributions of this paper are as follows:
  • We design a frequency-aware loss using DWT to sharpen fine details during client updates, and introduce a complementary aggregation method to weight model contributions more effectively. To the best of our knowledge, this is the first algorithm specifically designed to address VSR in the FL setting, effectively bridging the gap between privacy preservation and reconstruction quality.
  • Our approach makes no architecture-specific assumptions and does not require storing past weights, making it stateless and model-agnostic. This design reduces memory usage, avoids privacy risks, and eliminates synchronization overhead in federated learning on edge devices.
@@
  • Extensive experiments validate FedVSR’s effectiveness, outperforming existing approaches and paving the way for privacy-preserving low-level vision in the FL setting.
@@ 3.1 VSR Problem Definition
In addition to the widely used RGB color space, other representations such as the YUV and YCbCr color spaces are frequently employed in VSR applications. Let $F_{i}\in\mathbb{R}^{H\times W\times 3}$ denote the $i$-th frame in a given LR video sequence, while $\hat{F}_{i}\in\mathbb{R}^{sH\times sW\times 3}$ represents its corresponding high-resolution (HR) counterpart, where $s$ is the scaling factor (e.g., $s=2,4,8$). Additionally, we define a set of $12N+$ 1 HR frames $\{\hat{I}_{j}\}_{j=i-N}^{i+N}$ centered around $\hat{I}_{i}$, where $N$ represents the temporal radius. The degradation process that generates an LR video sequence from its HR counterpart can be mathematically formulated as:
@@
A more detailed expression for the degradation process can be written as:
@@
where $D$ represents the down-sampling operation, $B$ denotes the blurring effect, $E_{i\to j}$ is a warping function that accounts for motion between frames $\hat{I}_{i}$ and $\hat{I}_{j}$, and $n_{j}$ represents the image noise. The goal of a VSR algorithm is to reconstruct an HR video sequence from degraded LR frames in a manner that closely resembles the ground truth video. This makes this task an ill-posed problem because there exist infinitely many possible HR images $\hat{I}_{i}$ that can produce the same LR image $I_{i}$ after degrading, meaning the degradation operator is highly non-injective [49]. While some of the techniques used in ISIR can be applied to individual frames, VSR must also consider inter-frame dependencies to ensure motion coherence across the sequence. The inverse process of degradation, which represents the super-resolution task, can be expressed as:
@@ 3.2 FL Problem Definition
FL aims to minimize the following objective function across participating clients:
@@
Each client updates the global model locally:
@@
where $\eta$ is the learning rate. The updated models are aggregated by the central server using a weighted averaging scheme:
@@ 4.1.1 DWT for High-Frequency Preservation
The coefficient in those four sub-bands are calculated as:
@@
$L_{i}^{LH}=\bar{F}_{i}^{LH}-\hat{F}_{i}^{LH}$ (9)
$L_{i}^{HL}=\bar{F}_{i}^{HL}-\hat{F}_{i}^{HL}$ (10)
$L_{i}^{HH}=\bar{F}_{i}^{HH}-\hat{F}_{i}^{HH}$ (11)
Then with the Charbonnier loss [9] we calculate the helper term $\mathcal{L}_{\text{DWT}}$ :
@@
Now, we define the overall loss for the model with two weighted terms, one for the generic VSR loss $\mathcal{L}_{\text{VSR}}$ and one for the DWT-based high-frequency loss $\mathcal{L}_{\text{HiFr}}$ as follows:
@@ 5 Experiments
Model Algorithm Less Heterogeneous More Heterogeneous
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
VRT [30] FedAvg[36] 29.71 0.8522 0.1916 24.88 0.7680 0.2446 28.85 0.8282 0.2200 24.60 0.7516 0.2557
FedProx[27] 28.82 0.8281 0.2133 24.18 0.7341 0.2668 28.64 0.8236 0.2222 24.10 0.7256 0.2713
SCAFFOLD[21] 28.59 0.8181 0.2228 23.99 0.7182 0.2737 28.17 0.8074 0.2318 24.04 0.7208 0.2739
FedVSR (Ours) 29.86 0.8562 0.1908 24.92 0.7682 0.2393 29.03 0.8339 0.2123 24.73 0.7584 0.2492
RVRT [29] FedAvg[36] 31.30 0.8877 0.1536 25.96 0.8156 0.1974 29.89 0.8578 0.1867 25.67 0.8040 0.2065
FedProx[27] 30.09 0.8618 0.1899 25.20 0.7775 0.2330 29.59 0.8526 0.1943 25.02 0.7730 0.2353
SCAFFOLD[21] 29.63 0.7967 0.2101 24.79 0.7317 0.2407 26.71 0.7142 0.2590 23.08 0.6441 0.3018
FedVSR (Ours) 31.54 0.8922 0.1455 26.05 0.8191 0.1889 30.90 0.8811 0.1576 25.90 0.8131 0.1910
IART [54] FedAvg[36] 30.56 0.8729 0.1695 25.43 0.7949 0.2118 29.52 0.8493 0.1909 25.04 0.7774 0.2272
FedProx[27] 29.65 0.8500 0.1683 24.72 0.7599 0.2307 29.30 0.8444 0.1958 24.42 0.7505 0.2293
SCAFFOLD[21] 30.29 0.8678 0.1692 25.18 0.7857 0.2084 29.26 0.8413 0.1992 24.85 0.7690 0.2278
FedVSR (Ours) 30.88 0.8812 0.1556 25.54 0.8012 0.1990 29.76 0.8536 0.1887 25.26 0.7850 0.2220
Table 1: PSNR ($\uparrow$), SSIM ($\uparrow$), and LPIPS ($\downarrow$) comparison of FL algorithms across different VSR models and datasets under varying levels of heterogeneity.
@@
For our baseline, we use a combination of FedAvg[36], FedProx[27], and SCAFFOLD[21] from the FL side, along with VRT[30], RVRT[29], and IART[54] from VSR. We selected these FL methods as they are widely accepted and frequently used to benchmark core algorithmic performance in recent literature. While newer approaches like MOON[24], FedDyn[1], FedDisco[57], FedLAW[28] exist, they are rely on task-specific assumptions such as logits, class-wise statistics, or server-side data proxies that do not translate well to low-level, regression-based tasks like VSR. On the other hand, exisitng FL-SR models such as FedSR, p²FedSR, and FedMRI are not directly comparable to FedVSR for two key reasons. First, they are tightly coupled to a specific model architecture: they require dividing the model into predefined blocks (“shallow–deep–upsampler”), which limits flexibility and demands structural changes to existing SR models. Second, these methods are stateful meaning each client must store additional data across training rounds, such as degradation descriptors, deep feature extractors, or previous model weights. This leads to significantly higher storage, computation, and coordination overhead. In contrast, FedVSR is designed to be both model-agnostic and stateless, making it simpler, more lightweight, and easier to apply in diverse real-world scenarios.
@@
The evaluation metrics used for the experiments are Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS).
@@ 5.2 Results and Comparison
As shown in Table 1, our proposed FedVSR algorithm consistently outperforms other federated learning (FL) algorithms across all tested models and datasets. On average, FedVSR surpasses existing FL methods by a margin of 0.61 dB in PSNR and 0.022 in SSIM and 0.022 in LPIPS in less heterogeneous settings, while in more heterogeneous settings, the margins are 0.66 dB in PSNR and 0.028 in SSIM and 0.019 in LPIPS, demonstrating its robustness under challenging conditions. Among the evaluated models, RVRT benefits the most from FedVSR, achieving the highest performance margin. In contrast, IART shows the lowest overall improvement but still maintains a competitive edge over the baseline FL methods.
The observed performance trends in our experiments align with existing findings on FedProx and SCAFFOLD, further validating our results. As shown in Fig.2, FedVSR consistently achieves superior PSNR values across different VSR models and datasets, outperforming FedAvg, FedProx, and SCAFFOLD, particularly in later training rounds. This aligns with prior research on FedProx, which highlights its failure to reach the stationary points of the global optimization objective, even under homogeneous data distributions. The study suggests that FedProx’s limiting points may remain far from the true minimizer of the global empirical risk [48], which could explain its suboptimal performance in our experiments. Although FedProx may still generalize well by aggregating local data effectively in certain conditions, its weaker convergence properties become apparent in more heterogeneous scenarios. We also observe that FedProx is highly sensitive to the $\mu$ value during training. The original paper suggests $\{0.001,0.01,0.1,1\}$ as candidate values and finds that $\mu=1$ performs better in most cases. However, we notice that for any $\mu$ greater than 0.001, the VSR model struggles significantly with learning. This suggests that the relationship is more complex and requires further investigation and analysis.
@@
As showen in Fig. 1 clearly demonstrate the limitations of FedAvg, FedProx, and SCAFFOLD in accurately reconstructing fine details. While these methods reduce some pixelation from the LQ input, they all fail to recover the textures and sharp edges present in the ground-truth. FedAvg produces overly smooth results, blurring important structural details such as the wooden lattice and frame edges. FedProx exhibits similar blurring, struggling to restore high-frequency textures. SCAFFOLD, not only fails to produce a sharp reconstruction but also introduces noticeable artifacts, distorting the texture of the wooden lattice and making the output appear unnatural. In contrast, FedVSR delivers significantly sharper and more realistic reconstructions, clearly outperforming the other approaches. Due to space limitations, additional qualitative and quantitative results are provided in the supplementary material 8.